# Inferring Resource-Oriented Intentions using LLMs for Static Resource Leak Detection

### v3

@@ VI Discussion
VI-A Effects of Different LLMs
In our current implementation, InferROI utilizes the powerful LLM, GPT-4, which has demonstrated its effectiveness in inferring resource-oriented intentions based on the designed prompt template. It is important to consider the potential impact of using different LLMs with varying model parameter sizes and training dataset scales. For instance, LLMs such as GPT-3.5 might exhibit different performance levels with GPT-4. To gain insights into this aspect, we conducted some preliminary experiments with GPT-3.5 to qualitatively assess its effectiveness in resource-oriented intention inference. Our observations indicate that GPT-3.5 exhibits a notable performance gap, with approximately 20% lower precision and recall compared to GPT-4. This difference could be attributed to the more powerful analytical capabilities of GPT-4 in code understanding, as highlighted in recent investigation [9, 10]. These findings motivate further exploration of developing specific LLMs tailored for resource-oriented intention inference, potentially involving fine-tuning of “smaller” models like Llama [29, 30]. Moreover, although we set the temperature of the LLM as zero to remove randomness, the LLM outputs are still sometimes unstable. Therefore, we further repeat our experiments on a small set of cases and find the variance in the final detection performance is negligible.
VI-B Static Analysis
InferROI adopts a lightweight intraprocedural static analysis for detecting resource leaks. Instead of utilizing interprocedural analysis, we opt for an intraprocedural analysis based on the observation that many lower-level resource APIs (e.g., AndroidHttpClient) are encapsulated within higher-level resource APIs (e.g., Socket). By directly inferring intentions for higher-level resources, we can reduce the complexity of interprocedural program analysis during the detection process.
@@ VI-B Static Analysis
Nevertheless, it is important to acknowledge the limitations in the employed static analysis, especially in scenarios where certain types of resource leaks might remain undetected. For instance, resource leaks occurring within Android callbacks represent a challenging area where our lightweight static analysis may fall short. Our key insights lie in directly inferring resource-oriented intentions based on code context understanding to boost static leak detection. The advantage of inferring resource-oriented intentions is its potential applicability to various static leak detection techniques, including callback-aware analyses like Relda2 [17]. By integrating the inferred resource-oriented intentions into such existing techniques, we can enhance their capability to handle more comprehensive cases. In this regard, our LLM-based resource-oriented intention inference can provide complementary insights and improvement for existing static resource leak detection, with the potential to empower a wide range of analysis methods to achieve more comprehensive and accurate results.
VI-C Threats to Validity
@@
The internal validity of our studies is potentially affected by the randomness in data sampling and the subjectiveness in data annotation. To mitigate these threats, we adopted commonly-used data sampling strategies and involved multiple annotators to minimize any preference bias. Another concern is the risk of data leakage, where the evaluation data might have been incorporated into the training data of the utilized LLM. However, the ablation study (RQ3.b) already shows that purely querying GPT-4 itself to detect resource leak has poor effectiveness, indicating that the effectiveness of InferROI is mainly credited to our proposed intention inference and static analysis algorithm not to the data leakage. Furthermore, we further curate projects created after the date of the GPT-4’s training data collection (in RQ2) to mitigate this issue. As for the external validity, it is essential to consider the evaluation data used in our work, which might not guarantee the full generality of our findings. To address this concern, we evaluate InferROI on both existing datasets and real-world open-source projects to detect resource leaks.
@@ VII Related Work
Different from the existing detection approaches, our work does not rely on prepared API pairs and directly infers resource-oriented intention in code by utilizing the exceptional code comprehension capability of large language models.
LLMs for Software Engineering. Recent studies have demonstrated that pre-trained language models can be utilized as knowledge bases by using synthetic tasks similar to the pre-training objective to retrieve the knowledge/information stored in the models [35, 36]. These works have shown that language models can recall factual knowledge without any fine-tuning by using proper prompts. Several works have focused on exploring effective prompt templates to improve performance of PLMs on downstream problems. In a recent survey by Liu et al. [37], prompt templates are broadly classified into two categories: cloze prompts [38, 35], which entail filling in the blanks in text or code, and prefix prompts [39, 40], which continue generating content following a specified prefix. Recent studies have explored the application of prompt learning in various software engineering tasks. Wang et al. [41] investigated the effectiveness of prompt learning in code intelligence tasks such as clone detection and code summarization. Huang et al. [42] used prompt learning for type inference.
In our work, we consider LLMs as knowledge bases that contain resource management knowledge with powerful code understanding capability, and prompt LLMs to infer resource-oriented intentions for resource leak detection.
@@ VIII Conclusions and Future Work
In this work, we propose InferROI, a novel resource leak detection approach which boosts static analysis via LLM-based resource-oriented intention inference. Given a code snippet, InferROI prompts the LLM in inferring involved intentions from the code. By aggregating these inferred intentions, InferROI utilizes a lightweight static-analysis based algorithm to analyze control-flow paths extracted from the code, thereby detecting resource leaks. We evaluate InferROI for the Java programming language and investigate its effectiveness in resource leak detection. Experimental results on the DroidLeaks and JLeaks datasets demonstrate that InferROI exhibits promising bug detection rate (about 60%) and false alarm rate (about 20%). Moreover, when applied to open-source projects, InferROI identifies 26 unknown resource leak bugs. In the future, we will explore the extension of InferROI to other programming languages and integrate InferROI with more program analysis techniques.

### v4

@@ VI Discussion
VI-A Limitations
Intention Inference. Some reported false alarms arise because InferROI fails to recognize certain resource release intentions. For example, five false alarms in DroidLeaks result from InferROI not inferring that the Cursor resource is released when calling the Android method CursorAdapter.changeCursor(Cursor). This oversight may be due to the API name changeCursor not explicitly indicating its resource release functionality, causing InferROI to overlook it. In fact, as shown in Figure 5, the Android documentation specifies that changeCursor(Cursor) releases the existing Cursor resource. However, it appears that the LLM does not retain this information. To address this issue, future work can explore integrating retrieval-augmented generation (RAG) techniques to search relevant API documentation and achieve more accurate and comprehensive intention inference.
Figure 5: API Description of CursorAdapter.changeCursor(Cursor)
@@ VI-A Limitations
Static Analysis. It is important to acknowledge the limitations in the employed static analysis, especially in scenarios where certain types of resource leaks might remain undetected. For instance, resource leaks occurring within Android callbacks represent a challenging area where our lightweight static analysis may fall short. Our key insights lie in directly inferring resource-oriented intentions based on code context understanding to boost static leak detection. The advantage of inferring resource-oriented intentions is its potential applicability to various static leak detection techniques, including callback-aware analyses like Relda2 [17]. By integrating the inferred resource-oriented intentions into such existing techniques, we can enhance their capability to handle more comprehensive cases. In this regard, our LLM-based resource-oriented intention inference can provide complementary insights and improvement for existing static resource leak detection, with the potential to empower a wide range of analysis methods to achieve more comprehensive and accurate results.
VI-B Threats to Validity
@@
The internal validity of our studies is potentially affected by the randomness in data sampling and the subjectiveness in data annotation. To mitigate these threats, we adopted commonly-used data sampling strategies and involved multiple annotators to minimize any preference bias. Another concern is the risk of data leakage, as the evaluation data might have been incorporated into the training data of the utilized LLMs. To the best of our knowledge, there is no existing resource leak benchmark that is guaranteed to be exempted from data leakage issues (i.e., only containing data after GPT-4 training data cutoff). Therefore, to mitigate this issue, we have made the following attempts to justify that the improvements of InferROI are not simply caused by the potential data memorization. (i) We evaluate InferROI to find 26 previously-unknown bugs when scanning the open-source projects in RQ2. Even if the code itself could be seen by LLMs during its training, the previously-unknown bugs are not seen by LLMs before. (ii) We show InferROI outperforms basic GPT-4 in the ablation study (RQ3.b). It indicates that the improvements of InferROI do not simply come from the potential memorization of GPT-4 on similar examples from training. (iii) We include multiple datasets in RQ1 to address the overfitting issues. In addition to the classic benchmark DroidLeak, we also include the latest resource leak benchmark JLeaks which has shown to have better diversity and quality than previous benchmarks. As for the external validity, the evaluation data used in our work might not guarantee the full generality of our findings. To address this concern, we evaluate InferROI on both existing datasets and real-world open-source projects to detect resource leaks.
@@ VII Related Work
Our work does not rely on prepared API pairs and directly infers resource-oriented intentions in code by utilizing the exceptional code comprehension capability of LLMs.
LLMs for Software Engineering. LLMs have been widely used in software engineering tasks [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Recent studies have demonstrated that pre-trained language models can be utilized as knowledge bases by using synthetic tasks similar to the pre-training objective to retrieve the knowledge/information stored in the models [54, 55]. These works have shown that language models can recall factual knowledge without any fine-tuning by using proper prompts. Several works have focused on exploring effective prompt templates to improve performance of PLMs on downstream problems. In a recent survey by Liu et al. [56], prompt templates are broadly classified into two categories: cloze prompts [57, 54, 58], which entail filling in the blanks in text or code, and prefix prompts [59, 60], which continue generating content following a specified prefix. Recent studies have explored the application of prompt learning in various software engineering tasks. Wang et al. [61] investigated the effectiveness of prompt learning in code intelligence tasks such as clone detection and code summarization. Huang et al. [62] used prompt learning for type inference.
This works consider LLMs as knowledge bases of resource management and prompts them to infer resource-oriented intentions based on their powerful code understanding capability.
@@ VIII Conclusions and Future Work
In this work, we propose InferROI, a novel resource leak detection approach which boosts static analysis via LLM-based resource-oriented intention inference. Given a code snippet, InferROI prompts the LLM in inferring involved intentions from the code. By aggregating these inferred intentions, InferROI utilizes a lightweight static-analysis based algorithm to analyze control-flow paths extracted from the code, thereby detecting resource leaks. We evaluate InferROI for the Java programming language and investigate its effectiveness in resource leak detection. Experimental results on the DroidLeaks and JLeaks datasets demonstrate that InferROI exhibits promising bug detection rate (about 60%) and false alarm rate (about 20%). Moreover, when applied to open-source projects, InferROI identifies 29 unknown resource leak bugs. In the future, we will explore the extension of InferROI to other programming languages and integrate InferROI with more program analysis techniques.
Data Availability
To facilitate further research, we have released our code and data in the replication package [33].
Acknowledgement
This research project is supported by National Key R&D Program of China (2023YFB4503805) and National Natural Science Foundation of China under Grant No. 62302099.